OpenStudy (anonymous):

For f(x)=4x+1 and g(x)=x^2-5 find (f-g)(x). A) -x^2+4x-4 B)4x^2-19 C)-x^2+4x+6 D)x^2-4x-6

OpenStudy (solomonzelman):

basically, subtract what f(x) is equal to, from g(x) in this case, (g-f)(x) means (x^2-5)-(4x+1) can you calculate that?

OpenStudy (mertsj):

@SolomonZelman (f-g)(x) is what the problem says.

OpenStudy (solomonzelman):

Maybe I interpreted it the wrong way; sorry if I did.

OpenStudy (solomonzelman):

@KeishaCx , do you know how to calculate \[(x^2-5)~~-~~(4x+1)\]

OpenStudy (mertsj):

(f-g)(x)=4x+1-(x^2-5)=4x+1-x^2+5=-x^2+4x+6

OpenStudy (solomonzelman):

oh, yeah, my bad, the other way, sorry, Mertsj, continue.